dskoyr:

A circle is centered at the point (5, -4) and passes through the point (-3, 2). The equation of this circle is (x + __)2 + (y + __)2 = ? mod note: this question is asking to complete the equation of the circle

justus:

What is your question?

Nnesha:

\[\rm (x-h)^2+(y-k)^2=r^2\] is the equation for a circle where (h,k) is the center and "r" is the radius. replace (h,k) with the center point given and (x,y) with (-3,2) to find r.

Nnesha:

`A circle is centered at the point (5, -4)` `and passes through the point (-3, 2).` `The equation of this circle is (x + __)2 + (y + __)2 = ?` \[ (x+__)^2+(y+__)^2=?\]